Database last updated on Mon, 02 Feb 2026 22:10:53 +0000 by localhost
Instrument constant changes for the week before 2022307 (3 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 6 rows)
Measurements and tests summary.
| _jday | _loc | _serial | _skc |
| 307 | STONY PLAIN | #195 | ecsc |
(0 sec)
Dead time (0 sec)
Nothing to report.
Standard lamp R6 (2 sec)
Nothing to report.
Standard lamp R5 (1 sec)
Nothing to report.
Number of days since last dome/window cleaning (0 sec, 2 rows)
| serial | days |
| 195 | more than a year |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 195 | 7 | 0 | 3 | 0 | 0 | 0 | 0 | 0 | 0 |
Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (1 sec)
No ozone measurements to display.
Ozone results for good FM obs (std<5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)
No ozone measurements to display.
Unsorted comments (0 sec, 9648 rows)
| Serial | type | Text | Count | YYYYJJJ |
| 195 | env: | DosBox | 1 | 2022307 |
| 195 | cubdsp: | 1 | 1 | 2022307 |
| 195 | fre_after: | re 2670 2681 | 1 | 2022307 |
| 195 | fre_after: | re 2575 2586 | 1 | 2022307 |
| 195 | fre_after: | re 2326 2337 | 1 | 2022307 |
| 195 | fre_after: | re 2327 2338 | 2 | 2022307 |
| 195 | fre_after: | re 2569 2580 | 2 | 2022307 |
| 195 | re: | Motors initializing- | 2 | 2022307 |
| 195 | fre_after: | re 2571 2582 | 2 | 2022307 |
| 195 | re: | Motors initializing-> | 5 | 2022307 |
| 195 | re: | BREWER SPECTROPHOTOMETER #00195 KIPP | 6 | 2022307 |
| 195 | fre_after: | re 2669 2680 | 10 | 2022307 |
| 195 | fre_after: | re 2572 2583 | 19 | 2022307 |
| 195 | re: | Motors initializing | 25 | 2022307 |
| 195 | pd: | Must do HG now | 77 | 2022307 |
| 195 | hg: | Automatically adding HP | 77 | 2022307 |
| 195 | hp: | Attempting recovery RESET | 77 | 2022307 |
The END (0 sec, 2 rows)